OpenStudy (anonymous):

John had the average score of 85.5 on his first four Checkpoints. The first score was 68. The second score was 5.5 more than the fourth score. The third score was 1.5 less than the fourth score. What are his scores in the second, third and fourth Checkpoints? Show your work, not just final answer.

OpenStudy (anonymous):

85.5=(68+(5.5+x)+(x-1.5)+x)/4 342=68+(5.5+x)+(x-1.5)+x 274=5.5+x+x-1.5+x 274+4+3x 270=3x x=90

OpenStudy (anonymous):

so test 1=68 test2=95.5 test3=88.5 test4=90

OpenStudy (anonymous):

test=checkpoint

OpenStudy (anonymous):

wow...that looks really complicated. Thanks for your help :)

OpenStudy (anonymous):

your welcome